OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(2-1)/(-2-5)=-1/7

OpenStudy (alexandervonhumboldt2):

y=-1/7x+b

OpenStudy (alexandervonhumboldt2):

1=-1/7x+b

OpenStudy (alexandervonhumboldt2):

b=8/7 y=-1/7x+8/7

OpenStudy (alexandervonhumboldt2):

1/7x+y=8/7

OpenStudy (anonymous):

 this is how i need it to be , how do i put it like that?